OpenStudy (anonymous):

help please someone find the indicated outputs for f(x) = 5xto second power - 2x f(0) = --? f(-1) = --? f(2) = --?

OpenStudy (saifoo.khan):

f(0) = --? 0

OpenStudy (saifoo.khan):

f(-1) = --? 5

OpenStudy (saifoo.khan):

f(2) = --? 5/16

OpenStudy (anonymous):

broo.... teachhh dont give the fishiess ;)

OpenStudy (asadkarim7):

f( 0 ) = 0 f ( -1) = 5x^2 - 2x f ( -1) = 5 ( -1) ^2 - 2( -1) = 5 + 2 = 7 f(2) = 5 ( 2 )^ 2 -2( 2) = 20 - 4 = 16

OpenStudy (saifoo.khan):

Wtf does tht mean Arman?

OpenStudy (anonymous):

hi asadkarim so you have completed your answers